# Anyon Quasilocalization in a Quasicrystalline Toric Code

### v1

@@ I Introduction
Earlier work by Levine and others PhysRevB.34.596; PhysRevB.34.617; PhysRevB.32.5547; PhysRevB.37.8145; RevModPhys.63.699 has provided comprehensive classification and construction methods for quasicrystals. Being neither perfectly periodic nor completely disordered, quasicrystalline non-interacting tight-binding models exhibit rich spectral properties in both one and higher dimensions: they feature fully localized states as well as Cantor-like critical spectra that are neither localized nor extended PhysRevB.35.1020; PhysRevLett.56.2740; PhysRevB.38.1621. These spectral features lead to unusual dynamical and transport properties in both non-interacting PhysRevB.46.13751; PhysRevB.62.15569; PhysRevB.99.224204 and interacting systems lev2017transport; PhysRevB.100.104204; PhysRevB.106.184209; PhysRevB.106.094206.
@@
While strongly correlated Landau-like phases in quasicrystals have been explored, the study of topological phases in such systems remains relatively sparse. Recent works have examined free-fermionic symmetry-protected topological phases PhysRevLett.109.106402; PhysRevLett.123.196401; PhysRevLett.124.036803; PhysRevResearch.2.033071; PhysRevLett.129.056403; fan2022topological, whereas the study of topologically ordered phases is still in its early stages PhysRevB.101.115413; PhysRevX.11.041051. The latter—characterized by fractionalized excitations and long-range entanglement—are of central importance from both theoretical RevModPhys.89.041004; savary2016quantum; broholm2020quantum and quantum computational perspectives semeghini2021probing; satzinger2021realizing.
@@
Our study begins with the tri-coordinated quasicrystalline (QC) spin liquid reported in Ref. PhysRevB.110.214438. By taking the strong-$z$ anisotropy limit, we demonstrate that, independent of the lattice generation (Gen), the resulting toric code (TC) model exhibits a hierarchical structure of star and plaquette operators of various orders, accompanied by a hierarchy of exponentially separated energy scales. Notably, when this same methodology is applied to the honeycomb lattice, it yields a TC model that is translationally symmetric both in terms of operator content and coupling constants. We show that the effective QC toric code (QCTC) Hamiltonian possesses an intricate structure governed by the hierarchy of energy scales, which are intimately connected to the underlying QC geometry. Furthermore, the introduction of a perturbative magnetic field gives rise to qualitatively new phenomena absent in crystalline counterparts. In particular, states at the same energy can exhibit both localized and extended characteristics, with their dispersion under increasing field strength $h$ displaying distinct scaling behaviors depending on the nature of the corresponding wavefunctions. Remarkably, anyonic charges in the QCTC are found to be either fully localized or undergo a stepwise delocalization process, where the charge spreads sequentially over subsets of sites of increasing size. We refer to this behavior as quasilocalization, akin to the phenomenon identified by Passaro et al. in their seminal work on electron delocalization in quasicrystals under strong coupling conditions PhysRevB.46.13751. In addition, we find that beyond the energy hierarchy, the flux configurations in the ground state themselves induce interference effects that can localize states. These localized modes are immune to any perturbative magnetic field, arising purely from the interplay between the quasicrystal geometry and the intrinsic flux distribution.
@@ II Model
The coupling parameters $\lambda^{(n)}_{A/B}$ of $A_{v}$ and $B_{p}$ operators are obtained from the perturbation theory calculations (see SM Supp-note). For the $A_{v}$ terms, only the couplings corresponding to $n=2,3,4,5$ are nonzero, and to the leading order perturbation-note these are given by, $\lambda^{(2)}_{A}=+J^{2}/2J_{z}$, $\lambda^{(3)}_{A}=-3J^{3}/8J_{z}^{2}$, $\lambda^{(4)}_{A}=+J^{4}/16J_{z}^{3}$, and $\lambda^{(5)}_{A}=-35J^{5}/2^{7}J_{z}^{4}$. For $B_{p}$ operators, the non-vanishing coupling parameters are: $\lambda^{(4)}_{B}=+J^{4}/16J_{z}^{3}$, $\lambda_{B}^{(6)}=-3J^{6}/2^{8}J^{5}_{z}$, and $\lambda_{B}^{(10)}=-3J^{10}/2^{16}J^{9}_{z}$. Here, we have chosen $J_{x}=J_{y}=J$, see SM for the general expressions. All the operators present in the Hamiltonian commute with each other, which makes the model exactly solvable in absence of any perturbations. The model also preserves the time-reversal (TR) invariance of $H_{K}$ (Eq.(1)) note-time-rev.
@@ IV Action of Zeeman Fields
subjected to the flux constraints Eq.(12). In the following, our focus will be on understanding the low-field single-particle spectrum and localization dynamics of this effective model.
@@ V Single-particle Spectrum
Hereafter, we drop the subscript of the Zeeman field and denote it by $h$. Since $H_{b}$ is quadratic, the properties of its low-energy eigenstates can be easily determined from exact diagonalization of various generations. The results are summarized in Fig.3 and described as follows.
The energy spectrum of the model on the generation-5 lattice (251 sites), plotted as a function of the perturbative field $h$, reveals four well-separated energy bands centered around $E\sim 2|\lambda_{A}^{(2)}|g^{n}$, with $g=J/J_{z}$ and $n=0,1,2,3$ (Fig.3(a)) spectrum-note. These bands correspond to single $e$-anyon excitations residing on TC lattice sites with different coordination numbers, each associated with the flipping of $A_{v}^{(n)}$ operators of different orders (see Fig. 3(b)). Owing to the spatial non-uniformity in the onsite potentials (or couplings of $A_{v}^{(n)}$), the energy dispersion with $h$ exhibits both linear ($\sim h$) and quadratic ($\sim h^{2}$) scaling, as illustrated for selected low-lying states from the $A^{(5)}$ sector in Fig. 3(c) uniform-potential-comment.
@@ VIII Conclusion
While our analysis has focused on the TC phase in a specific quasicrystalline geometry, it will be equally interesting to look at this phase in other quasicrystalline geometries and especially in the amorphous setting to see how the chiral spin liquid transits to the gapped TC phase cassella2023exact; KimPRL2023. Additionally, phason fluctuations PhysRevB.34.3345; PhysRevLett.134.136003; PhysRevB.85.094434; PhysRevB.79.172406—local rearrangements of the tiling—may have nontrivial effects on the low-energy dynamics of anyons in the quasicrystal.
Although we have qualitatively argued for the nature of the fractionalized phase beyond perturbative Zeeman couplings and sketched the broad structure of the phase diagram (see Fig. 2), a complete quantitative understanding of the full phase diagram will require more computationally intensive methods, particularly to uncover how confinement physics develops in this settingPhysRevB.82.085114; PhysRevLett.117.210401; PhysRevX.11.041008; 3lxj-dx76; PhysRevB.105.075132; PhysRevResearch.6.013298. Even at the perturbative-field level, the anyon dynamics already exhibits features reminiscent of a Bose-glass yu2012bose, and it would be valuable to obtain a more quantitative characterization of such glassy phases within the context of topologically ordered $\mathbb{Z}_{2}$ spin-liquids PhysRevLett.114.247207; PhysRevLett.118.087203; zhu2025emergent; PhysRevB.108.165118; hart2021correlation. Similar glassy behavior has been reported previously in disordered toric code PhysRevB.83.075124. In translationally invariant systems, the field-induced confinement transition is typically understood as arising from the condensation of bosonic $e$ and $m$ charges. Here, however, geometric constraints prevent the formation of system-spanning delocalized states, raising the possibility that the confinement transition may not be second order. Moreover, the quasi-periodic distribution of background $\pi$-fluxes and the discrete rotational symmetries of the quasicrystal may play a key role near the transition—for instance, the critical theory (if the transition is second-order) could exhibit emergent symmetries borla2025oddtoriccode not apparent in the $\mathbb{Z}_{2}$ symmetric microscopic model.
@@ IX Acknowledgments
We acknowledge fruitful discussions with Subrata Pachhal. SS acknowledges support from Institute Postdoctoral Fellowship, IIT Kanpur. AA and MS acknowledge prior collaboration with Sunghoon Kim, Animesh Nanda, Debanjan Chowdhury and Subhro Bhattacharjee on related themes. MS also acknowledges financial support from Indian Institute of Technology Kanpur as well as from the Department of Physics, University of Illinois Urbana-Champaign. AA acknowledges support from IITK Grants (IITK/PHY/2022010) and (IITK/PHY/2022011). Numerical calculations were performed on the workstations Wigner and Syahi at IITK.
@@ II Derivation of quasicrystalline toric code model
Fourth order: In this order of expansion, one $A_{v}$ and one $B_{p}$ term is generated. The $A_{v}$ ($B_{p}$) term arises from octagons (squares) having only $z$ and $y$ ($x$ and $y$) bonds at the boundaries. Calculation of $A_{v}^{(4)}$ follows exactly the same steps as done for $A_{v}^{(3)}$ and we obtain,
$H_{\text{eff},1}^{(4)}=-\frac{J_{y}^{4}}{16J_{z}^{3}}A_{v}^{(4)}\ \ \text{where},\ \ A_{v}^{(4)}=\prod_{j=1}^{4}\tau^{x}_{j}$ (S12)
@@
Fifth order: In this order, $B_{p}$ terms do not arise, which is consistent with the time-reversal symmetry. A non-zero $A_{v}$ contribution arises from the decagonal plaquette (of the tri-coordinated lattice). There are $5!$ possible orderings of the $yy$ perturbations. The possible sequences of the intermediate state energy denominators are: ($0\rightarrow 4\rightarrow 4\rightarrow 4\rightarrow 4\rightarrow 0$), ($0\rightarrow 4\rightarrow 4\rightarrow 8\rightarrow 4\rightarrow 0$), ($0\rightarrow 4\rightarrow 8\rightarrow 4\rightarrow 4\rightarrow 0$), and ($0\rightarrow 4\rightarrow 8\rightarrow 8\rightarrow 4\rightarrow 0$). After collecting all of these processes, we obtain the following numerical pre-factor
$5\times(i^{2})^{5}\times\frac{(-J_{y})^{5}}{J_{z}^{4}}\bigg[\frac{8}{(-4)^{4}}+\frac{8}{(-4)^{2}(-8)^{2}}+\frac{8}{(-4)^{3}(-8)}\bigg]$ (S15)
and
$H_{\text{eff}}^{(5)}=+\frac{35}{128}\frac{J_{y}^{5}}{J_{z}^{4}}A_{v}^{(5)}\ \ \text{where}\ A_{v}^{(5)}=\prod_{j=1}^{5}\tau^{x}_{j}$ (S16)
@@
Tenth order: The next non-trivial corrections come at this order of expansion. Again, we neglect the “disconnected plaquette” contributions, for example $\sim(A^{(m)}_{v})_{i}(B^{(n)}_{p})_{j}$, with $m+n<10$, since these terms only trivially renormalize the dominant lower order coupling constants.
@@
$(i)\ \ W^{(4)}_{p}(x,z)=(-1)\prod_{r=1}^{4}\sigma^{y}_{r}=(-1)\prod_{j=1}^{2}\tau^{x}_{j}=-A_{s}^{(2)}\ \ \$$(ii)\ \ W^{(4)}_{p}(x,y)=(-1)\prod_{r=1}^{4}\sigma^{z}_{r}=(-1)\prod_{j=1}^{4}\tau^{z}_{r}=-B_{p}^{(4)}$
$(iii)\ \ W^{(6)}_{p}(x,z)=(-1)\prod_{r=1}^{6}\sigma^{y}_{r}=(+1)\prod_{j=1}^{3}\tau^{x}_{j}=A_{s}^{(3)}\ \ \$$(iv)\ \ W^{(6)}_{p}(y,z)=(-1)\prod_{r=1}^{6}\sigma^{x}_{r}=(-1)\prod_{j=1}^{3}\tau^{x}_{r}=-A_{s}^{(3)}$
$(v)\ \ W^{(6)}_{p}(x,y)=(-1)\prod_{r=1}^{6}\sigma^{z}_{r}=(-1)\prod_{j=1}^{6}\tau^{z}_{r}=-B_{p}^{(6)}\ \ \$$(vi)\ \ W^{(8)}_{p}(y,z)=(-1)\prod_{r=1}^{8}\sigma^{x}_{r}=(-1)\prod_{j=1}^{4}\tau^{x}_{r}=-A_{s}^{(4)}$
$(vii)\ \ W^{(10)}_{p}(y,z)=(-1)\prod_{r=1}^{10}\sigma^{x}_{r}=(-1)\prod_{j=1}^{5}\tau^{x}_{r}=-A_{s}^{(5)}\ \ \$$(viii)\ \ W^{(10)}_{p}(x,y)=(-1)\prod_{r=1}^{10}\sigma^{z}_{r}=(-1)\prod_{j=1}^{5}\tau^{z}_{r}=-B_{p}^{(10)}$ (S18)

### v2

@@ I Introduction
Earlier work by Levine and others PhysRevB.34.596; PhysRevB.34.617; PhysRevB.32.5547; PhysRevB.37.8145; RevModPhys.63.699 has provided comprehensive classification and construction methods for quasicrystals. Being neither perfectly periodic nor completely disordered, quasicrystalline non-interacting tight-binding models exhibit rich spectral properties in both one and higher dimensions: they feature fully localized states as well as Cantor-like critical spectra that are neither localized nor extended PhysRevB.35.1020; PhysRevLett.56.2740; PhysRevB.38.1621. These spectral features lead to unusual dynamical and transport properties in both non-interacting PhysRevB.46.13751; PhysRevB.62.15569; PhysRevB.99.224204 and interacting systems Vidal-Dominique-Giamarchi; lev2017transport; PhysRevB.100.104204; PhysRevB.106.184209; PhysRevB.106.094206.
@@
While strongly correlated Landau-like phases in quasicrystals have been explored, the study of topological phases in such systems remains relatively sparse. Recent works have examined free-fermionic symmetry-protected topological phases PhysRevLett.109.106402; Hofstadter-Fuchs-vidal; PhysRevLett.123.196401; PhysRevLett.124.036803; PhysRevResearch.2.033071; PhysRevLett.129.056403; fan2022topological, whereas the study of topologically ordered phases is still in its early stages PhysRevB.101.115413; PhysRevX.11.041051. The latter—characterized by fractionalized excitations and long-range entanglement—are of central importance from both theoretical RevModPhys.89.041004; savary2016quantum; broholm2020quantum and quantum computational perspectives semeghini2021probing; satzinger2021realizing.
@@
Our study begins with the tri-coordinated quasicrystalline (QC) spin liquid reported in Ref. PhysRevB.110.214438. By taking the strong-$z$ anisotropy limit, we demonstrate that, independent of the lattice generation (Gen), the resulting TC model exhibits a hierarchical structure of star and plaquette operators of various orders, accompanied by a hierarchy of exponentially separated energy scales. Notably, when this same methodology is applied to the honeycomb lattice, it yields a TC model that is translationally symmetric both in terms of operator content and coupling constants. We show that the effective QC toric code (QCTC) Hamiltonian possesses an intricate structure governed by the hierarchy of energy scales, which are intimately connected to the underlying QC geometry. Furthermore, the introduction of a perturbative magnetic field gives rise to qualitatively new phenomena absent in crystalline counterparts. In particular, states at the same energy can exhibit both localized and extended characteristics, with their dispersion under increasing field strength $h$ displaying distinct scaling behaviors depending on the nature of the corresponding wavefunctions. Remarkably, anyonic charges in the QCTC are found to be either fully localized or undergo a stepwise delocalization process, where the charge spreads sequentially over subsets of sites of increasing size. We refer to this behavior as quasilocalization, akin to the phenomenon identified by Passaro et al. in their seminal work on electron delocalization in quasicrystals under strong coupling conditions PhysRevB.46.13751. In addition, we find that beyond the energy hierarchy, the flux configurations in the ground state themselves induce interference effects that can localize states. These localized modes are immune to any perturbative magnetic field, arising purely from the interplay between the quasicrystal geometry and the intrinsic flux distribution.
@@ II Model
The coupling parameters $\lambda^{(n)}_{A/B}$ of $A_{v}$ and $B_{p}$ operators are obtained from the perturbation theory calculations (see SM Supp-note). For the $A_{v}$ terms, only the couplings corresponding to $n=2,3,4,5$ are nonzero, and to the leading order perturbation-note these are given by, $\lambda^{(2)}_{A}=+J^{2}/2J_{z}$, $\lambda^{(3)}_{A}=-3J^{3}/8J_{z}^{2}$, $\lambda^{(4)}_{A}=+5J^{4}/16J_{z}^{3}$, and $\lambda^{(5)}_{A}=-35J^{5}/2^{7}J_{z}^{4}$. For $B_{p}$ operators, the non-vanishing coupling parameters are: $\lambda^{(4)}_{B}=+J^{4}/16J_{z}^{3}$, $\lambda_{B}^{(6)}=-3J^{6}/2^{8}J^{5}_{z}$, and $\lambda_{B}^{(10)}=-3J^{10}/2^{16}J^{9}_{z}$. Here, we have chosen $J_{x}=J_{y}=J$, see SM for the general expressions. All the operators present in the Hamiltonian commute with each other, which makes the model exactly solvable in absence of any perturbations. The model also preserves the time-reversal (TR) invariance of $H_{K}$ (Eq.(1)) note-time-rev.
@@ IV Action of Zeeman Fields
subjected to the flux constraints Eq.(12). The transverse field $\lambda_{v}^{(n)}$ depends on a vertex $v$ solely through its coordination number $n$, and $\lambda_{v}^{(n)}=\lambda_{A}^{(n)}$. In the following, our focus will be on understanding the low-field single-particle spectrum and localization dynamics of this effective model.
@@ V Single-particle Spectrum
Hereafter, we drop the subscript of the Zeeman field and denote it by $h$. From now on, we also ignore the precise numerical factors present in $\lambda_{v}^{(n)}$ (or $\lambda_{A}^{(n)}$), and take $|\lambda_{v}^{(n)}|=J_{z}g^{n}$, where $g=J/J_{z}$, and $n=2,3,4,5$. Since $H_{b}$ is quadratic, the properties of its low-energy eigenstates can be easily determined from exact diagonalization of various generations. The results are summarized in Fig.3 and described as follows.
The energy spectrum of the model on the generation-5 lattice (251 sites), plotted as a function of the perturbative field $h$, reveals four well-separated energy bands centered around $E\sim 2|\lambda_{A}^{(2)}|g^{n}$ and $n=0,1,2,3$ (Fig.3(a)) spectrum-note. These bands correspond to single $e$-anyon excitations residing on TC lattice sites with different coordination numbers, each associated with the flipping of $A_{v}^{(n)}$ operators of different orders (see Fig. 3(b)). Owing to the spatial non-uniformity in the onsite potentials (or couplings of $A_{v}^{(n)}$), the energy dispersion with $h$ exhibits both linear ($\sim h$) and quadratic ($\sim h^{2}$) scaling, as illustrated for selected low-lying states from the $A^{(5)}$ sector in Fig. 3(c) uniform-potential-comment.
@@ VIII Conclusion
While our analysis has focused on the TC phase in a specific quasicrystalline geometry, it will be equally interesting to look at this phase in other quasicrystalline PhysRevB.108.104208 and hyperbolic mx1t-74dm geometries, and especially in the amorphous setting to see how the chiral spin liquid transits to the gapped TC phase cassella2023exact; KimPRL2023. Additionally, phason fluctuations PhysRevB.34.3345; PhysRevLett.134.136003; PhysRevB.85.094434; PhysRevB.79.172406—local rearrangements of the tiling—may have nontrivial effects on the low-energy dynamics of anyons in the quasicrystal.
Although we have qualitatively argued for the nature of the fractionalized phase beyond perturbative Zeeman couplings and sketched the broad structure of the phase diagram (see Fig. 2), a complete quantitative understanding of the full phase diagram will require more computationally intensive methods, particularly to uncover how confinement physics develops in this settingPhysRevB.82.085114; PhysRevLett.117.210401; PhysRevX.11.041008; 3lxj-dx76; PhysRevB.105.075132; PhysRevResearch.6.013298. Even at the perturbative-field level, the anyon dynamics already exhibits features reminiscent of a Bose-glass yu2012bose, and it would be valuable to obtain a more quantitative characterization of such glassy phases within the context of topologically ordered $\mathbb{Z}_{2}$ spin-liquids PhysRevLett.114.247207; PhysRevLett.118.087203; zhu2025emergent; PhysRevB.108.165118; hart2021correlation. Similar glassy behavior has been reported previously in disordered TC PhysRevB.83.075124. In translationally invariant systems, the field-induced confinement transition is typically understood as arising from the condensation of bosonic $e$ and $m$ charges. Here, however, geometric constraints prevent the formation of system-spanning delocalized states, raising the possibility that the confinement transition may not be second order. Moreover, the quasi-periodic distribution of background $\pi$-fluxes and the discrete rotational symmetries of the quasicrystal may play a key role near the transition—for instance, the critical theory (if the transition is second-order) could exhibit emergent symmetries borla2025oddtoriccode not apparent in the $\mathbb{Z}_{2}$ symmetric microscopic model.
@@ IX Acknowledgments
We acknowledge fruitful discussions with Subrata Pachhal. SS acknowledges support from Institute Postdoctoral Fellowship, IIT Kanpur. AA and MS acknowledge prior collaboration with Sunghoon Kim, Animesh Nanda, Debanjan Chowdhury and Subhro Bhattacharjee on related themes. We acknowledge Julien Vidal for discussions, critical reading of the manuscript and pointing out an error in an earlier version. MS also acknowledges financial support from Indian Institute of Technology Kanpur as well as from the Department of Physics, University of Illinois Urbana-Champaign. AA acknowledges support from IITK Grants (IITK/PHY/2022010) and (IITK/PHY/2022011). Numerical calculations were performed on the workstations Wigner and Syahi at IITK.
@@ II Derivation of quasicrystalline toric code model
Fourth order: In this order of expansion, one $A_{v}$ and one $B_{p}$ term is generated. The $A_{v}$ ($B_{p}$) term arises from octagons (squares) having only $z$ and $y$ ($x$ and $y$) bonds at the boundaries. Calculation of $A_{v}^{(4)}$ follows exactly the same steps as done for $A_{v}^{(3)}$. There are $4!$ equivalent orderings of $V$ and two different sequence of intermediate energy denominators: ($0\rightarrow 4\rightarrow 4\rightarrow 4\rightarrow 0$) and ($0\rightarrow 4\rightarrow 8\rightarrow 4\rightarrow 0$). After adding all of thenm, we obtain,
$H_{\text{eff},1}^{(4)}=-\bigg[\frac{16}{64}+\frac{8}{128}\bigg]\frac{J_{y}^{4}}{J_{z}^{3}}A_{v}^{(4)}=-\frac{5J^{4}_{y}}{16J^{3}_{z}}A^{(4)}_{v}\ \ \text{where},\ \ A_{v}^{(4)}=\prod_{j=1}^{4}\tau^{x}_{j}$ (S12)
@@
Fifth order: In this order, $B_{p}$ terms do not arise, which is consistent with the time-reversal symmetry. A non-zero $A_{v}$ contribution arises from the decagonal plaquettes (of the tri-coordinated lattice). There are two types of them: (1) plaquettes that contain only $z$ and $y$ bonds, (2) plaquettes that contain only $z$ and $x$ bonds (see Fig.S3). There are $5!$ equivalent orderings of $yy$ and $xx$ perturbations. The possible sequences of the intermediate state energy denominators (in both cases) are: ($0\rightarrow 4\rightarrow 4\rightarrow 4\rightarrow 4\rightarrow 0$), ($0\rightarrow 4\rightarrow 4\rightarrow 8\rightarrow 4\rightarrow 0$), ($0\rightarrow 4\rightarrow 8\rightarrow 4\rightarrow 4\rightarrow 0$), and ($0\rightarrow 4\rightarrow 8\rightarrow 8\rightarrow 4\rightarrow 0$). After collecting all of these processes, we obtain the following numerical pre-factor
$\text{For plaquette type-(1): }\ \ \ \ 5\times(i^{2})^{5}\times\frac{(-J_{y})^{5}}{J_{z}^{4}}\bigg[\frac{8}{(-4)^{4}}+\frac{8}{(-4)^{2}(-8)^{2}}+\frac{8}{(-4)^{3}(-8)}\bigg]$ (S15)
$\text{For plaquette type-(2): }\ \ \ \ 5\times\frac{(-J_{x})^{5}}{J_{z}^{4}}\bigg[\frac{8}{(-4)^{4}}+\frac{8}{(-4)^{2}(-8)^{2}}+\frac{8}{(-4)^{3}(-8)}\bigg]$ (S16)
Therefore,
$H_{\text{eff},1}^{(5)}=+\frac{35}{128}\frac{J_{y}^{5}}{J_{z}^{4}}A_{v}^{(5)}\ \ ,\ \ H_{\text{eff},2}^{(5)}=-\frac{35}{128}\frac{J_{x}^{5}}{J_{z}^{4}}A_{v}^{(5)}\ \ ,\text{where}\ A_{v}^{(5)}=\prod_{j=1}^{5}\tau^{x}_{j}$ (S17)
@@
Tenth order: The next non-trivial correction comes at this order of expansion, which generates $B_{p}^{(10)}$. There are total 10! possible ways of acting the perturbations and we numerically determine the corresponding coupling constant,
$H_{\text{eff}}^{(10)}=+\frac{35}{2^{16}}\frac{J_{x}^{5}J_{y}^{5}}{J_{z}^{9}}B_{p}^{(10)},\ \ \ \text{where }\ B_{p}^{(10)}=\prod_{j=1}^{10}\tau^{z}_{j}$ (S19)
Here also, we neglect the “disconnected plaquette” contributions, for example $\sim(A^{(m)}_{v})_{i}(B^{(n)}_{p})_{j}$, with $m+n<10$, since these terms only trivially renormalize the dominant lower order coupling constants.
@@
$(1)\ \ W^{(4)}_{p}(x,z)=(-1)\prod_{r=1}^{4}\sigma^{y}_{r}=(-1)\prod_{j=1}^{2}\tau^{x}_{j}=-A_{v}^{(2)}\ \ \$$(2)\ \ W^{(4)}_{p}(x,y)=(-1)\prod_{r=1}^{4}\sigma^{z}_{r}=(-1)\prod_{j=1}^{4}\tau^{z}_{r}=-B_{p}^{(4)}$
$(3)\ \ W^{(6)}_{p}(x,z)=(-1)\prod_{r=1}^{6}\sigma^{y}_{r}=(+1)\prod_{j=1}^{3}\tau^{x}_{j}=A_{v}^{(3)}\ \ \$$(4)\ \ W^{(6)}_{p}(y,z)=(-1)\prod_{r=1}^{6}\sigma^{x}_{r}=(-1)\prod_{j=1}^{3}\tau^{x}_{r}=-A_{v}^{(3)}$
$(5)\ \ W^{(6)}_{p}(x,y)=(-1)\prod_{r=1}^{6}\sigma^{z}_{r}=(-1)\prod_{j=1}^{6}\tau^{z}_{r}=-B_{p}^{(6)}\ \ \$$(6)\ \ W^{(8)}_{p}(y,z)=(-1)\prod_{r=1}^{8}\sigma^{x}_{r}=(-1)\prod_{j=1}^{4}\tau^{x}_{r}=-A_{v}^{(4)}$
$(7)\ \ W^{(10)}_{p}(y,z)=(-1)\prod_{r=1}^{10}\sigma^{x}_{r}=(-1)\prod_{j=1}^{5}\tau^{x}_{r}=-A_{v}^{(5)}\ \ \$$(8)\ \ W^{(10)}_{p}(x,z)=(-1)\prod_{r=1}^{10}\sigma^{y}_{r}=(+1)\prod_{j=1}^{5}\tau^{x}_{r}=A_{v}^{(5)}$
$(9)\ \ W^{(10)}_{p}(x,y)=(-1)\prod_{r=1}^{10}\sigma^{z}_{r}=(-1)\prod_{j=1}^{5}\tau^{z}_{r}=-B_{p}^{(10)}$ (S20)